OpenStudy (troy121101):

Which statement correctly describes the relationship between the graph of f(x) and g(x)=f(x−3)? The graph of g(x) is the graph of ​f(x) translated 3 units down. The graph of g(x) is the graph of ​f(x)​ translated 3 units left. The graph of g(x) is the graph of ​f(x)​ translated 3 units up. The graph of g(x) is the graph of ​f(x) translated 3 units right.

OpenStudy (royalranger):

The graph moves about the axis. That narrows it down.

OpenStudy (royalranger):

x axis*

OpenStudy (troy121101):

so left or right?

OpenStudy (troy121101):

eliminate answer choice A and C? so it is either B or D?

OpenStudy (royalranger):



satellite73 (satellite73):

\[g(x)=f(x−3)\] say for example \[f(x)=x^2, g(x)=f(x-3)=(x-3)^2\]

satellite73 (satellite73):

then the graphs look identical, but the first one has vertex \((0,0)\) and the second one has vertex \((3,0)\) os the graph has shifted RIGHT

OpenStudy (troy121101):

oh, I see, so D would be correct, sorry, I didn't get much out of @royalranger 's chart, thank you for clearing that